### v2

# Why Train Everything? Tint a Single Layer for Multi-task Model Merging

Abstract
Model merging integrates independently fine-tuned models into a single multi-task model, offering a flexible alternative to joint training. However, many existing model merging methods introduce additional task-specific components, increasing complexity and requiring extra modifications. We propose Model Tinting , a lightweight yet highly effective approach that improves model merging by updating just a single layer, accounting for as low as 0.5% of total parameters. Our key observation is that explicit task-specific modules are not necessary; instead, subtle adjustments to a single layer can effectively capture task-specific variations within the merged model while maintaining generalization. We introduce a confidence-based filtering mechanism to alleviate the impact of unreliable predictions from individual models on the merged model. Extensive experiments across vision and NLP tasks demonstrate that Model Tinting achieves state-of-the-art performance, even in challenging dense prediction tasks. Our code is available at https://github.com/AIM-SKKU/ModelTinting.
1 Introduction
Multi-task learning (MTL) [4] enables models to handle multiple tasks efficiently by facilitating knowledge transfer. Model merging [67, 22, 68] emerged as an alternative to conventional MTL [76, 40, 28, 20] by combining independently fine-tuned models into a single model. This approach allows for flexible adaptation, reduces computational overhead, and better preserves task-specific knowledge. A key challenge in model merging is balancing generalization and task specialization. On one hand, a merged model should retain shared knowledge across tasks [72, 74], but on the other hand, it must preserve task-specific adaptations that are crucial for individual task performance [73, 21].
Figure 1: Performance comparison of model merging methods. Radar charts highlight Model Tinting’s superiority across (a) 20 vision tasks and (b) 8 NLP tasks, achieving state-of-the-art results.
Early methods [23, 72, 75, 74, 71] primarily focused on constructing task-agnostic models by averaging weights or aligning parameters across models, assuming that similar models in weight space would naturally form a meaningful shared representation. However, this generalization-first strategy often led to a loss of task-specific information, reducing performance on individual tasks. This issue became more pronounced when merging models trained on diverse or conflicting tasks, as simple averaging often failed to preserve specialized adaptations essential for high accuracy.
Recent approaches [21, 73, 58, 36] explored task-specific parameterization, with some incorporating additional components, in some cases, are trainable to enhance adaptation. These methods typically added task-specific modules or learnable merging coefficients, selectively adjusting how different tasks influence the final model. While this enhanced task adaptation, it also introduced challenges such as increased model complexity, higher computational costs in both training and inference, and scalability issues when applied to many tasks. Furthermore, many merging strategies still relied on a fixed shared model [23, 72, 75, 11, 36, 21, 73, 58], often formed by merging encoders without adaptation. As a result, the merged model struggled to retain fine-grained task distinctions and failed to effectively leverage task-relevant cues, leading to suboptimal performance in diverse task distributions (see Figure 1). These challenges raise two fundamental questions: 1) Can task-specific features be captured without introducing additional modules? 2) Is a predefined task-agnostic backbone sufficient to retain task-specific knowledge effectively?
In this paper, we propose Model Tinting , a simple yet effective model merging method that preserves task-specific knowledge with minimal overhead. Instead of introducing additional modules, our method maintains task-specific information through minimal trainable adjustments, updating a single layer that accounts for fewer than 0.5% of parameters in ViT-B/32. By jointly optimizing task-specific layers and merging coefficients, it refines both components under the guidance of individually fine-tuned models. To further enhance reliability, we introduce a confidence-based filtering mechanism that prevents the model from inheriting incorrect predictions from individual experts. By selectively removing unreliable supervision signals, this mechanism improves robustness and enables the merged model to sometimes surpass individual task models.
Pub. Method Task-specific Adaptation No Extra Module Adaptive Shared Model
ICLR 2023 Task Arithmetic [23]
ICLR 2024 AdaMerging [74]
ICML 2024 Surgery [73]
ICML 2024 WEMoE [58]
NeurIPS 2024 EMR-Merging [21]
NeurIPS 2024 TWIN-Merging [36]
Model Tinting
Table 1: Comparison of recent model merging methods.
Table 1 shows the main differences between existing methods and our method. Model Tinting requires no additional modules beyond a single trainable layer per task, making it lightweight and easy to deploy. Furthermore, we also adjust the task-agnostic model (i.e., shared model) to better leverage the effect of the task-specific layer, further enhancing adaptation for each task. Through extensive experiments across image classification, dense prediction, and NLP tasks, we demonstrate that Model Tinting achieves state-of-the-art performance across different model scales and task configurations. Notably, our method excels in dense prediction tasks, a challenging yet underexplored domain in prior model merging studies. Our approach achieves remarkably strong results, demonstrating its ability to handle complex task structures effectively. Our main contributions are as follows:
  • We propose a lightweight MTL model merging method that preserves task-specific knowledge with minimal trainable adjustments, requiring no additional modules.
  • To improve robustness, we introduce a confidence-based filtering mechanism that selectively removes unreliable supervision signals from individual models, preventing error propagation in the merged model.
  • Our method achieves state-of-the-art performance across diverse tasks, including image classification, dense prediction, and NLP, particularly excelling in dense prediction, an underexplored area in model merging.
2 Related Work
2.1 Multi-task Learning
Multi-task Learning (MTL) enables models to perform multiple tasks simultaneously by leveraging shared knowledge [60, 4, 78]. However, MTL faces challenges like task interference and negative transfer. To address these issues, modularization techniques introduce task-specific modules or pathways, preserving unique task information and reducing interference [37, 40, 50]. Gradient-based methods balance tasks through normalization and align gradient directions to minimize conflicts [6, 76]. Task weighting approaches dynamically assign weights to balance the learning process and prevent any single task from dominating [28, 51, 34, 20, 17]. Knowledge distillation further enhances MTL by transferring insights from specialized models to a unified framework [13, 25, 70]. Despite these advancements, traditional MTL methods often require extensive labeled data and significant computational resources. This highlights the need for more efficient and scalable approaches such as model merging, which can consolidate independently fine-tuned models into a unified framework.
2.2 Model Merging for Multi-task Learning
Negative transfer remains a fundamental challenge in model merging for MTL, as merging often inherits task interference within the shared representation. The goal is to construct a single model that generalizes well across all tasks while minimizing such interference. Existing weight alignment methods [71, 55, 1] enable merging fine-tuned weights obtained from different pre-trained models. However, these approaches face scalability limitations as the number of tasks increases. A foundational work in model merging for MTL is Task Arithmetic [23], which introduces task vectors, defined as the difference between pre-trained and fine-tuned models. This enables arithmetic-based weight combinations across multiple tasks, allowing the merging of a large number of individual models. Subsequent methods refine this approach by explicitly addressing task conflict within the merged model. Ties-Merging [72] mitigates redundancy and sign conflicts, while DARE [75] removes redundant components and rescales key parameters. AdaMerging [74] learns task vector importance at test time, and LiNeS [64] reduces the impact of shallow layers to preserve general representations.
While merged models have traditionally been considered task-agnostic, recent findings [63, 19] suggest that they inherently encode task-specific knowledge. Motivated by these observations, recent approaches incorporate task-specific modules into the merged encoder to better align with the performance of individual models. Surgery [73] refines merged encoders by introducing task-specific adapters, while EMR-Merging [21] applies task-specific masking to a pre-defined merged model. Router-based methods [58, 36] determine task vector coefficients via routing mechanisms that depend on each task’s input. However, these methods introduce additional parameters and often overlook the role of a well-structured shared expert in maximizing task-specific module effectiveness. Model Tinting explicitly refines both shared and task-specific layers in a complementary manner, which allows it to easily capture hidden task-specific knowledge. This balanced design improves task adaptation and often surpasses individual model performance.
3 Background
3.1 Preliminary
Problem definition. Multi-task Learning (MTL) [4] aims to train a single model to achieve strong performance across multiple tasks. Given the input $x^{k}\in\mathbb{R}^{d_{in}}$ for the kth task, a $L$-layer encoder $f(x^{k};\phi)$ with the weights $\phi=\{\phi^{1},\ldots,\phi^{L}\}$ outputs the representation $\hat{z}^{k}\in\mathbb{R}^{d_{out}}$. For each of the $K$ tasks, $\hat{z}^{k}$ is fed into a task-specific layer $g(\cdot\ ;\theta_{k})$, where $\theta_{k}$ represents weights unique to task $k$, to perform classification (e.g., the classification head in CLIP [48]). We employ two types of weights to address MTL: $\phi_{\text{pre}}$ and $\phi_{k}$, which denote a pre-trained weight and task-specific fine-tuned weights for task $k$, respectively. These weights form the basis of novel weight merging methods [67, 26] aimed at excelling in MTL.
Task Arithmetic [23] introduced task vectors for each task $k$, defined as $\tau_{k}=\phi_{k}-\phi_{\text{pre}}$, to average weights for MTL. The merged weight is represented by $\phi_{\text{MTL}}=\phi_{\text{pre}}+\lambda\cdot\sum_{k=1}^{K}\tau_{k}$, where $\lambda$ is a scaling hyperparameter for the task vectors, determined using held-out validation sets rather than being learned. With the merged weight and individual classifiers, a single merged model $f(x;\phi_{\text{MTL}})$ is capable of managing all tasks with decent performance.
AdaMerging [74] improves Task Arithmetic by constructing a merged model $f(x;\phi_{\text{MTL}})$ through learning $\lambda$ in a task- and layer-wise manner, defined as $\phi_{\text{MTL}}^{l}=\phi_{\text{pre}}^{l}+\sum_{k=1}^{K}\lambda_{k}^{l}\cdot\tau_{k}^{l}$. Each merging coefficient $\{\lambda_{k}^{1},\ldots,\lambda_{k}^{L}\}$ is trained by optimizing: $\underset{\{\lambda_{k}^{l}\}}{\min}\sum_{k,x^{k}\in\mathcal{X}}\mathcal{L}_{k}\left(g_{k}\left(f(x^{k};\phi_{\text{MTL}});\theta_{k}\right)\right).$ Without ground-truth labels (for test set $\mathcal{X}$), AdaMerging adopts an entropy minimization objective $\mathcal{L}_{k}$ to learn each coefficient.
Figure 2: Target and cross-task evaluations across 8 tasks using ViT-B/32 with base and trained classifiers. We also evaluate both Task Arithmetic and AdaMerging encoders. Training classifiers on merged representations improves cross-task generalization while preserving target task performance. Additionally, performance varies depending on the shared encoder’s ability to retain task-specific features.
3.2 Pilot Study
The success of recent MTL merging methods [36, 21, 58, 73] lies in introducing additional task-specific components to extract hidden task-specific knowledge [19, 63] from a predefined merged encoder (i.e., shared encoder) [^1] Those MTL merging methods generally achieve overall improvements across tasks through weight averaging, which suggests that the shared encoder becomes more likely to be task-agnostic.. Instead, we argue that selecting appropriate layers within the model is sufficient without introducing additional modules. Furthermore, we suggest that it is crucial to adjust the shared encoder to better capture task-specific information.
Study outline. Our goal is to validate this claim and extend it further by demonstrating that a trained task-specific layer can enhance task-specific capabilities even within a task-agnostic representation. Additionally, we investigate whether the shared encoder affects the extraction of task-specific knowledge. We use the classifier as our task-specific layer for analysis due to its simplicity. If our hypothesis holds, merging methods [74, 73] could achieve improved performance with minimal additional training.
Task-specific training. We generate task-agnostic representations from the fixed merged encoder $f(x^{k};\phi_{\text{MTL}})$, which is obtained using Task Arithmetic [23] or AdaMerging [74]. Then, we train a base classifier $g_{k}(\cdot;\theta_{k})$ from the pre-trained model for the corresponding $k^{th}$ task (e.g., $k=1,...,8$). The following objective is minimized with respect to the weights ($\theta_{k}^{tr}$) of the selected layer ($g_{k}^{tr}$) during training:
$\underset{\{\theta_{k}^{tr}\}}{\min}\sum_{k,(x^{k},y^{k})\in\mathcal{X}}\mathcal{L}_{k}(g_{k}^{tr}(f(x^{k};\phi_{\text{MTL}});\theta_{k}^{tr}),y^{k}),$ (1)
where $g_{k}^{tr}$ represents the trainable classifier and $y^{k}$ denotes the corresponding ground-truth label for the input sample $x^{k}$. This setup uses labeled samples to confirm the encoder’s maximal capability by utilizing ground-truth labels.
Target & cross-task evaluation. We evaluate each base $g_{k}^{bs}$ and trained classifier $g_{k}^{tr}$ using features from all individual encoders $f(\cdot;\phi_{\text{m}})$ (e.g., $648^{2}=$ 64 cases). To assess its effectiveness, we compute the average performance in two settings: 1) target task evaluation, where the classifier is tested on features from its corresponding encoder (i.e., $m=k$), and 2) cross-task evaluation, where it is tested on features from individual encoders trained on distinct tasks (i.e., $m\neq{}k$).
Interpretation. Figure 2 shows the average performance of target task and cross-task evaluations separately, using base classifiers and trained classifiers. We observe consistent improvements in cross-task performance when using the classifiers trained on other tasks. Detailed performance for each task pair is provided in Appendix C.1. These results suggest that training the classifier on merged features also facilitates the alignment of individual features with other tasks. This demonstrates the trained layer’s enhanced capability to refine task-specific outputs from task-agnostic representations. We further argue that an intermediate layer in the backbone could function similarly, and this will also be empirically proven by our experiments (see Figure 4).
Additionally, the choice of the shared encoder used for classifier training also affects performance. Interestingly, we observe that when using the Task Arithmetic features, cross-task performance improves significantly more than with the AdaMerging features. This highlights the importance of identifying a suitable shared encoder that better leverages task-specific knowledge. A remaining challenge is that weight merging is typically unavailable to access the ground-truth labels in practice. Therefore, our method is designed to handle this constraint and remains effective in real-world scenarios.
4 Method
4.1 Model Tinting
Motivation. Our pilot study revealed that training a single layer (e.g., a classifier head) on merged representations enhances task-specific performance and can improve overall MTL effectiveness. Building on this finding, we propose an efficient merging method that fine-tunes a selected single layer in the merged encoder for each task, extending beyond standard adaptation approaches, while also refining the merged model.
Task-specific adaptation at test time. Our method follows test-time adaptation strategies [74, 73], where ground-truth labels are unavailable. A common approach is to use entropy [74] as a proxy for supervision. However, entropy only roughly follows the trend of cross-entropy [74, 44] and may not align well when data cardinality is insufficient. Our empirical analysis (see Table 2) shows that entropy does not consistently approximate cross-entropy with ground-truth supervision during training. Thus, we propose a more effective way to incorporate cross-entropy.
Building on our pilot study (refer to Section 3.2), we align the classifier’s output with each task encoder’s knowledge to inject task-specific information. We define an objective to align each fine-tuned model’s distribution with that of the merged model on the test set $\mathcal{X}$. Formally, for each task $k$, we minimize the discrepancy between $C_{k}^{\text{ft}}=\{g_{k}^{\text{bs}}(f(x^{k};\phi_{k});\theta_{k}^{\text{bs}})\mid x^{k}\in\mathcal{X}\}$ and $C_{k}^{\text{merged}}=\{g_{k}^{\text{tr}}(f(x^{k};\phi_{\text{MTL}});\theta_{k}^{\text{tr}})\mid x^{k}\in\mathcal{X}\}$.
Instead of relying on semi-supervised techniques [77, 59, 46, 2], we adopt a simpler approach [33]: training the merged model to match the fixed output distributions of the individual models [^2] While individual models could be updated during training or replaced with other merged models as alternative options, we opt for a simpler approach. We provide the related study in Figure 5. using the cross-entropy loss. This optimization objective is defined as follows:
$\underset{\{\lambda_{k}^{l}\},\{\theta_{k}^{\text{tr}}\}}{\min}\;\sum_{k=1}^{K}\mathcal{L}_{k}\left(C_{k}^{\text{merged}},C_{k}^{\text{ft}}\right),$ (2)
where $\mathcal{L}_{k}$ represents the loss for each task, $g_{k}^{\text{bs}}$ are the base task-specific layers, and $g_{k}^{\text{tr}}$ are the learnable task-specific layers. Unlike AdaMerging [74], which optimizes only merging coefficients in the shared encoder, our approach jointly optimizes both coefficients and task-specific layers to better tailor representations for distinct tasks.
SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD
Entropy Before 0.66 0.72 0.80 0.86 0.92 0.92 1.00 0.56
After 0.32 -0.47 0.79 0.94 0.83 0.41 0.96 0.47
$\Delta$ (-0.34) (-1.19) (-0.01) (+0.08) (-0.09) (-0.51) (-0.04) (-0.09)
Ours Before 0.80 0.86 0.97 1.00 0.98 0.99 1.00 0.84
After 0.82 0.88 0.99 1.00 0.99 1.00 1.00 0.88
$\Delta$ (+0.02) (+0.02) (+0.02) (+0.00) (+0.01) (+0.01) (+0.00) (+0.04)
Table 2: Spearman correlation of losses with ground truth cross-entropy loss. We report the correlation coefficient for two merged weights: ‘before’ (initial merged weights) and ‘after’ (trained merged weights under the different losses). We evaluate two losses: ‘Entropy’ (the entropy minimization objective) and ‘Ours’ (our self-label-based cross-entropy loss).
Loss correlation. To validate our approach, we analyze how well different loss functions correlate with ground-truth-based cross-entropy loss, which serves as a strong supervision signal when labels are available. We compute the Spearman correlation [^3] Spearman correlation quantifies monotonic relationships and is often used for ranked data or non-linear associations. between model-generated losses and the ground-truth-based cross-entropy loss. A higher correlation indicates that the loss function provides a reliable training signal, even in the absence of explicit labels.
Table 2 shows that entropy initially correlates moderately with supervised cross-entropy. However, after training, this correlation significantly weakens, and in some cases, entropy diverges sharply (e.g. Cars: -1.19 drop). This suggests that entropy is not a stable proxy for supervision and may fail to maintain consistency as training progresses. This finding aligns with prior studies [74, 44], which observed that entropy-based losses can misrepresent model confidence, leading to unreliable optimization. In contrast, our model-generated cross-entropy loss starts with a higher correlation and maintains strong alignment with supervised cross-entropy throughout training. Unlike entropy, which fluctuates unpredictably, our approach preserves a stable and reliable training signal across tasks.
4.2 Model Tinting++
Refining Model Tinting . Model-generated supervisory signals contribute to performance improvement, but not all signals are equally beneficial. In particular, cases where the merged model correctly classifies a sample while individual models fail may introduce unreliable training data. To address this, we assume that only high-confidence samples from individual models should be used for training. Prior studies [16] suggest that well-calibrated models mitigate overconfidence and achieve greater robustness. Following this principle, confidence-aware filtering is expected to enhance the stability of the merged model [44].
Directly setting a fixed confidence threshold is challenging and may introduce unwanted biases. Instead, we propose a relative confidence-based filtering strategy: We compare the top-1 probabilities of individual and merged models and exclude samples where the merged model assigns a higher probability than the individual models. This ensures that the merged model receives supervision primarily from reliable samples, leading to more stable adaptation. As discussed in [43], filtering uncertain cases enhances model robustness, which is expected to improve MTL performance across various tasks.
(a) Prediction discrepancy
(b) Overall difference
Figure 3: Prediction discrepancies between the merged model and individual model. (a) The upper bars indicate predictions correctly classified by individual models but misclassified by the merged model; the lower (hatched) bars indicate the opposite. A lower upper portion and a higher lower portion indicate a positive impact on the merged model performance. (b) represents the overall difference between the upper and lower bars.
Prediction discrepancy analysis. To evaluate the impact of our approach, we analyze discrepancies between merged model predictions and those of individual models, which often serve as a performance upper bound in MTL. Figure 3 presents two sets of bar graphs illustrating prediction inconsistencies. The upper bars represent samples correctly classified by individual models but misclassified by the merged model, while the lower hatched bars indicate cases where individual models fail but the merged model succeeds. A method with lower upper bars and higher lower bars is preferable, as it signifies reduced merging-induced errors and improved generalization.
Among the evaluated approaches, our method achieves the smallest prediction discrepancy, significantly reducing the number of misclassified samples. This suggests that the filtering strategy effectively retains the strengths of individual models while mitigating degradation caused by merging. Figure 3 further summarizes the overall impact by computing the net difference between the upper and lower bars. A smaller value in this metric indicates a more balanced model that preserves individual model accuracy while capturing additional gains. The proposed method consistently outperforms alternatives, demonstrating its effectiveness in minimizing errors while improving generalization, leading to enhanced MTL performance.
Method CoLA SST2 MRPC STSB QQP MNLI QNLI RTE Avg.
Individual 0.6018 0.9404 0.8922 0.9063 0.9141 0.8720 0.9271 0.7906 0.8556
[…]
Weight Averaging 0.1396 0.6411 0.6936 0.3184 0.7536 0.4219 0.5870 0.5523 0.5134
RegMean [27] 0.3667 0.9060 0.7574 0.6268 0.8355 0.7002 0.8235 0.5848 0.7001
Task Arithmetic [23] 0.1878 0.8589 0.7990 0.7403 0.8378 0.5908 0.6967 0.6209 0.6665
[…]
Ties-Merging [72] 0.2048 0.8440 0.8113 0.5819 0.8570 0.6465 0.7481 0.4296 0.6404
[…]
EMR-Merging [21] 0.3996 0.9335 0.8627 0.8277 0.8972 0.8545 0.8957 0.7437 0.8018
[…]
Model Tinting
0.6018 0.9392 0.8922 0.9049 0.8606 0.8382 0.8955 0.7906 0.8404
Table 3: Multi-task performance across 8 NLP tasks. We present a performance comparison of merging the RoBERTa models fine-tuned on 8 tasks in GLUE. $\dagger$ denotes our reproduced results for a fair comparison.
Method SUN Cars RES. Euro SVHN GTS. MNIST DTD Avg.
Pretrained 62.3 59.7 60.7 45.5 31.4 32.6 48.5 43.8 48.0
Individual 75.3 77.7 96.1 99.7 97.5 98.7 99.7 79.4 90.5
Traditional MTL 73.9 74.4 93.9 98.2 95.8 98.9 99.5 77.9 88.9
Weight Averaging 65.3 63.4 71.4 71.7 64.2 52.8 87.5 50.1 65.8
Fisher Merging [39] 68.6 69.2 70.7 66.4 72.9 51.1 87.9 59.9 68.3
RegMean [27] 65.3 63.5 75.6 78.6 78.1 67.4 93.7 52.0 71.8
Task Arithmetic [23] 55.2 54.9 66.7 78.9 80.2 69.7 97.3 50.4 69.1
[…]
Ties Merging [72] 65.0 64.4 74.8 77.4 81.2 69.3 96.5 54.5 72.9
[…]
Concrete TA [56] 62.5 61.1 76.0 95.7 91.0 81.9 98.5 51.9 77.3
Concrete AM [56] 67.8 70.0 87.5 96.0 91.6 96.7 98.7 63.8 84.0
AdaMerging++ [74] 66.6 68.3 82.2 94.2 89.6 89.0 98.3 60.6 81.1
Ada. w/ Surgery [73] 71.2 72.0 92.3 99.0 92.2 97.9 99.0 76.1 87.5
EMR-Merging [21] 75.2 72.8 93.5 99.5 96.9 98.1 99.6 74.4 88.7
WEMoE [58] 74.1 77.4 93.7 99.1 96.2 98.9 99.6 76.4 89.4
Model Tinting
73.7 77.6 94.7 99.0 95.4 98.2 99.1 79.1 89.6
Model Tinting++
74.3 79.3 94.8 99.0 95.7 98.5 99.2 80.2 90.1
(a) On merging ViT-B/32 models
Method SUN Cars RES. Euro SVHN GTS. MNIST DTD Avg.
Pretrained 66.8 77.7 71.0 59.9 58.4 50.5 76.3 55.3 64.5
Individual 82.3 92.4 97.4 100 98.1 99.2 99.7 84.1 94.2
Traditional MTL 80.8 90.6 96.3 96.3 97.6 99.1 99.6 84.4 93.5
Weight Averaging 72.1 81.6 82.6 91.9 78.2 70.7 97.1 62.8 79.6
Fisher Merging [39] 69.2 88.6 87.5 93.5 80.6 74.8 93.3 70.0 82.2
RegMean [27] 73.3 81.8 86.1 97.0 88.0 84.2 98.5 60.8 83.7
Task Arithmetic [23] 73.9 82.1 86.6 94.1 87.9 86.7 98.9 65.6 84.5
[…]
Ties Merging [72] 76.5 85.0 89.3 95.7 90.3 83.3 99.0 68.8 86.0
[…]
Concrete TA [56] 74.6 86.2 89.0 96.7 93.6 93.4 99.1 66.9 87.4
Concrete AM [56] 77.8 91.2 92.1 97.0 94.4 97.9 99.0 79.5 91.1
AdaMerging++ [74] 79.4 90.3 91.6 97.4 93.4 97.5 99.0 79.2 91.0
Ada. w/ Surgery [73] 80.3 90.8 94.3 98.2 94.1 98.7 99.2 82.5 92.3
EMR-Merging [21] 83.2 90.7 96.8 99.7 97.9 99.1 99.7 82.7 93.7
WEMoE [58] 81.4 92.6 95.4 99.4 97.7 99.3 99.7 83.7 93.6
Model Tinting
81.4 92.6 96.9 99.4 97.6 99.0 99.4 84.2 93.8
Model Tinting++
81.2 92.3 97.0 99.7 97.3 99.1 99.4 84.8 93.9
(b) On merging ViT-L/14 models
Table 4: Multi-task performance comparison across 8 vision tasks. We compare our methods with the competing methods for merging ViT-B/32 (left) and ViT-L/14 (right) fine-tuned models. $\dagger$ denotes reproduced results.
5 Experiment
5.1 Experimental Setup
Datasets and metrics. We follow the standard setups in DARE [75] and EMR-Merging [21] for NLP tasks, using fine-tuned weights for eight tasks from the GLUE benchmark [62]. Task-specific metrics are employed: Matthews correlation for CoLA [65], Pearson and Spearman correlations for STS-B [5], and accuracy for other tasks [53, 12, 24, 66, 49, 14]. For vision tasks, we adopt the setups from Task Arithmetic [23] and AdaMerging [74], using fine-tuned weights on eight image classification tasks. We extend our experiments to 12 additional tasks [32, 15, 42, 45, 61, 9, 10, 69, 3, 8, 47], covering a total of 20 tasks as in [63]. All vision tasks are evaluated using accuracy. Results for a subset of 14 tasks are provided in the Appendix C.3. For dense prediction evaluation, we use NYUv2 [52], which contains 1,449 RGB-D indoor scene images. It includes three tasks: 13-class semantic segmentation, depth estimation, and surface normal estimation. We evaluate them with mIoU/pixel accuracy, absolute/relative error, and mean angular error, respectively.
Models. For NLP tasks, we use RoBERTa-base [35], leveraging publicly available fine-tuned weights [^4] https://github.com/harveyhuang18/EMR_Merging from EMR-Merging [21]. We conduct experiments on TWIN-Merging [36] and our method under the same settings, using Mean Squared Error (MSE) loss for STS-B. For vision experiments, we employ pre-trained CLIP [48] models [^5] https://github.com/mlfoundations/open_clip, specifically ViT-B/32, B/16, L/14. Most analyses use ViT-B/32. To ensure comparability across eight tasks, we use publicly available fine-tuned weights [^6] https://github.com/mlfoundations/task_vectors from Task Arithmetic [23]. For 14- and 20-task experiments, we fine-tune CLIP ViT-B/32 following the Task Arithmetic configuration. For dense prediction, we adopt ResNet-50 [18] as the backbone, following FusionBench [57]. The model is initialized with ImageNet pre-trained weights and fine-tuned for each task. Since prior works lack results on 14- and 20-task settings, we reproduce multiple previous methods [38, 72, 64, 63, 74, 73, 21].
Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD CIFAR100 FER2013
[…]
Task $\dagger$ [23] 61.8 53.0 61.9 57.5 49.8 44.6 77.9 45.6 65.4 41.4
[…]
Model Tinting
72.9 76.1 94.5 98.7 92.4 97.4 99.1 79.5 84.7 69.2
Model Tinting++
73.5 78.6 94.5 99.2 93.3 97.2 99.3 79.7 83.9 68.5
Method Flowers Pet PCAM STL10 CIFAR10 EMNIST FMNIST Food101 KMNIST R-SST2
[…]
Task $\dagger$ 63.1 86.0 65.8 94.4 91.5 39.6 73.9 72.1 12.2 57.8
[…]
Model Tinting
90.9 91.9 82.1 97.8 96.0 98.3 90.6 85.5 95.3 75.2
Model Tinting++
92.0 92.4 84.7 98.0 96.0 98.8 91.2 85.2 96.3 76.4
Method Avg.
[…]
Task $\dagger$ 60.8
[…]
Model Tinting
88.4
Model Tinting++
88.9
Table 5: Task-specific and average performance across 20 vision tasks. We use ViT-B/32 for experiments. $\dagger$ denotes reproduced results.
Method Segmentation Depth Estimation Normal
[…]
Model Tinting
49.0 73.0 45.7 18.6 26.4
Model Tinting++
50.2 73.4 45.1 18.6 26.1
Table 6: Multi-task performance across three dense prediction tasks. We compare the performance of merging ResNet50 models fine-tuned on three dense prediction tasks in NYUv2. $\dagger$ denotes reproduced results.
5.2 Main Results
Merging 8 NLP tasks. As shown in Table 3, our Model Tinting largely improved over the best method on average. While EMR-Merging suffers from a significant performance drop on CoLA, and TWIN-Merging on STS-B, Model Tinting demonstrates consistent performance close to individual task-specific models across all tasks. We argue that this reveals the limitations of methods that rely on a fixed shared encoder and shows the superiority of our method, which effectively manages task-specific information from the adaptive shared encoder.
Merging 8 vision tasks. Table 4 and Table 4 show that our method outperforms traditional MTL methods, improving ViT-B/32 and ViT-L/14 by 0.7%p and 0.2%p over state-of-the-art methods (ViT-B/16 improves by 0.7%p; see Appendix C.3). Notably, for ViT-B/32, this is achieved with just 0.5M trainable parameters, far fewer than the 7M used in prior state-of-the-art approach. Additionally, our method achieves performance close to the individual models regardless of task difficulty and even surpasses them on certain tasks. For instance, on DTD, a particularly challenging dataset, other merging methods exhibit a noticeable performance gap compared to individual models, whereas our approach achieves a higher accuracy of 80.2% compared to 79.4%. When applying sample filtering, we observe a 0.5%p improvement for ViT-B/32, while the performance gain for ViT-L/14 is minimal. This is likely because Model Tinting already performs nearly as well as individual models when using ViT-L/14.
Merging 20 vision tasks. We explore the impact of scaling model merging by expanding the task set from 8 to 20, making the problem more complex. To ensure a fair comparison, we reproduce the performance of existing methods under the same experimental conditions. In Table 5, we observe that for KMNIST, both task arithmetic and AdaMerging exhibit notably low performance. Furthermore, on datasets like PCAM and EMNIST, AdaMerging occasionally results in lower performance than task arithmetic. In contrast, our approach shows a significant improvement, with a 19.3%p performance increase over AdaMerging and surpassing state-of-the-art methods by 2.3%p. Notably, for several tasks, our method even exceeds the performance of individual models. With sample selection, we see an additional 0.5%p performance boost, highlighting the effectiveness of our approach as the number of tasks scales.
Merging dense prediction tasks. Dense prediction tasks, such as segmentation, depth estimation, and normal estimation, pose significant challenges due to their inter-task heterogeneity. Despite this, most model merging studies focus on classification, leaving these tasks largely unexplored. Table 6 presents results for these tasks, highlighting the limitations of previous methods. Surgery, which performs comparably to individual models on eight classification tasks, suffers substantial performance drops in depth and normal estimation. Similarly, EMR-Merging, despite leveraging task-specific dynamic masks, struggles to maintain segmentation performance. In contrast, our approach effectively balances these tasks, maintaining performance close to individual models even in dense prediction settings.
5.3 Emprical Analyses
Effect of adjusting different layers. A natural question is whether the improvements observed in Model Tinting are specific to training the classifier. To explore this, we analyze the effect of training different layers beyond the classifier by updating merging coefficients along with each of the 12 transformer layers in the merged encoder. As shown in Figure 4, training a single internal layer achieves accuracy comparable to training the classifier, suggesting that task-specific information can be effectively captured by refining a single layer rather than modifying the entire model. Identifying the optimal layer could further enhance performance. On the other hand, training multiple layers at once, either from the early stage (layers 1–6) or the later stage (layers 7–12), leads to a significant drop in performance. This suggests that updating too many layers disrupts the task-agnostic knowledge embedded in the pre-trained model, reducing its ability to generalize across tasks.
Figure 4: Impact of training different task-specific layer. We compare the average performance of merged models when different task-specific layers are trainable, while other layers update only their merging coefficients. Training a single layer achieves accuracy similar to tuning the classifier, whereas modifying multiple layers leads to performance degradation.
Sparsity analysis of merging coefficients. Inspired by the literature [75, 72, 11], which highlights how redundant parameters degrade performance due to conflicts among task-specific parameters, we investigate this phenomenon in our method. We analyze the learned merging coefficients to observe their impact, specifically exploring sparsity after training them alone and jointly with the classifier. We find 37.2% versus 55.9% of the coefficients are concentrated near zero (i.e., smaller than 1e-5) with the deviated average accuracy of 83.4% versus 89.6% for each configuration. This suggests that training a task-specific layer, like the classifier, along with merging coefficients, reduces task interference by pruning unnecessary parameters, leading to improved accuracy. For visualizations of merging coefficients, please refer to Appendix D.4.
More studies with merging coefficients. We study employing individual model predictions without any refinement is a sensible choice for supervisory signals. We refine predictions by merging individual models with a pre-trained model, which is expected to yield more precise predictions. Specifically, we handle each individual model as a pre-trained model with an added task vector scaled by a merging coefficient, ranging from 0 (pre-trained model) to 1 (individual model). We experiment across these settings, and the results are shown in Figure 5, where performance consistently improved as predictions approached the individual model. This indicates that guiding the classifier with a task-agnostic pre-trained model limits its ability to capture task-specific details effectively. Though a combined model (at the coefficient 0.8) performs the best, which is about 89.9%, our design choice of using individual models alone achieves a near-best result. Finding the optimal coefficient would require cumbersome hyperparameter tuning, so using the existing individual models directly is a practical option. Efficiently searching for the coefficient would be a potential research direction in the future.
(a) On supervisory model
(b) Training robustness
Figure 5: More analyses with merging coefficients. (a) Refined supervisory models merged under various coefficients are tested. Our design choice – using the unmerged individual models – performs near-optimally. (b) MTL accuracy confirms the initialization robustness of our method across different initial merging coefficients; related results are illustrated in Figure 6.
Initialization robustness. Our method benefits from robust merging coefficient initialization. To highlight this, we present training curves under different initialization settings and compare them with those of other model merging methods. As shown in Figure 5, when only the classifier is trained (i.e., fixed shared encoder), the performance is sensitive to initialization; However, training both the classifier and coefficients together improves robustness. Figure 6 shows that when training a task-specific adapter only, the Surgery model is sensitive to initialization. Unlikely, our method jointly optimizes the merging coefficient and task-specific layer, making it robust to initialization sensitivity.
Figure 6: Impact of initializations. We test two different approaches: fixed-value initialization from Task Arithmetic (dashed lines) and 2) learned-value one in AdaMerging (solid lines). The learning trend diverges with what to learn, but training both coefficients and task-specific layers is not affected by the initializations.
6 Conclusion
In this paper, we have introduced Model Tinting , an efficient multi-task model merging approach that enhances task-specific knowledge with minimal overhead. By adaptively learning task-specific layers while refining the shared model, our method balances generalization and specialization without additional task-specific modules. This distinguishes our method from conventional model merging methods, which often introduce unnecessary complexity or struggle to retain fine-grained task information. To further improve robustness, Model Tinting++ incorporates a confidence-based filtering mechanism that selectively removes unreliable supervision signals. Extensive experiments across vision, NLP, and dense prediction tasks demonstrate that our approach consistently achieves state-of-the-art performance, particularly excelling in dense prediction and large-scale multi-task benchmarks. Additionally, our analysis reveals key insights: 1) Inner layers complement each other in capturing task-specific features; 2) Training task-specific layers induces sparsity in merging coefficients, improving task separation; 3) Our joint optimization strategy is robust to initialization, ensuring stable results across tasks. We believe this work provides a strong foundation for future research in scalable and adaptive model merging.
Appendix
This Appendix provides an overview of our experimental setup and baselines, further experimental results across tasks, and additional empirical analyses. The detailed descriptions of each section are summarized as follows:
  • Appendix A: Outline of Model Tinting with introducing an illustrative figure;
  • Appendix B: Detailed experimental setup and hyperparameter configurations;
  • Appendix C: Additional results for pilot study, NLP, and vision tasks;
  • Appendix D: Further analyses on task-specific layers, loss functions, convergence, and sparsity.
Appendix A Overview of Model Tinting
We give a brief overview of our proposed method Model Tinting , a test-time method that merges multiple fine-tuned models using a single trainable layer per task Figure A provides an illustrative overview of our method. At the top, the figure illustrates the merging process, where multiple fine-tuned models ($\phi_{1},\phi_{2},\ldots,\phi_{K}$) are combined into a shared representation ($\phi_{\text{MTL}}$) using adaptive merging coefficients. During this process, a trainable task-specific layer ($\phi_{\text{tint}}^{l}$) is introduced at a chosen transformer layer $l$, capturing task-specific information from the task-agnostic representations of the merged encoder. We argued that any layer can be picked for task adjustments before merging, and this is further supported by the graph in Figure A. The results show that applying Model Tinting, regardless of the chosen layer, leads to significant performance improvements over the baseline. This highlights the robustness and effectiveness of task-specific adjustments in enhancing task-agnostic representations, with more details provided in Appendix D.1. We also proposed Model Tinting++ , which applies a strategy to filter noisy predictions from individual models.
Appendix B Experimental Setup
B.1 Training Details
Main experiments. We initialize the layer-wise merging coefficients for all layers to 0.3 by default, following the values used in the previous method [74]. However, for the 14 and 20 vision tasks, the coefficients are initially set to 0.1, as using 0.3 incurred significantly lower Task Arithmetic [23] performance for these tasks, presumably due to the increased number of tasks. We use the Adam [29] optimizer with momentum parameters (0.9, 0.999) to update the coefficients and the classifier. For vision tasks, the learning rate is set to 0.01 when training the classifier and 0.001 when training only the merging coefficients. For dense prediction tasks, the learning rate of 0.0001 is used. For NLP tasks, the learning rate of 0.0005 is used, regardless of whether the classifier, the merging coefficients, or both are being updated. Consistent with prior works [74, 58, 73], vision tasks are trained with 16,000 test samples per task over 500 iterations, using a batch size of 32. Dense prediction tasks use 400 samples per task over 25 iterations with a batch size of 16. NLP tasks are trained with 32,000 samples per task over 1,000 iterations with a batch size of 32. For each task, the loss is computed, and the model parameters are updated immediately following the forward pass for each batch. To enable sequential processing, the order of tasks is randomized.
Figure A: Illustration of   Model Tinting . Our method improves multi-task learning by introducing a trainable, task-specific layer during model merging. Our approach (red) consistently achieves better results than the baseline without tinting (green) across all layers.
Dicussions on implementation. Previous methods [74, 58] sample a batch for each task, pass each batch through a shared backbone, and then process it through the corresponding task-specific head, which is typical. The losses from each task are then combined, and a single gradient update is performed after completing the forward pass for all tasks. However, this approach becomes less efficient as the number of tasks or model size increases.
To address this, we experiment with a sequential update strategy, where updates are applied immediately after each task’s forward pass instead of the typical one. Interestingly, this sequential update approach gives a positive byproduct: performance improvements over the traditional approach. While the original AdaMerging achieves an average accuracy of 80.1%, our sequential update strategy improves it to 81.6% for the 8 vision tasks using ViT-B/32. Similarly, our method also enjoys a gain when using sequential updates (89.6% compared to 89.4%). We argue that this improvement may result from mitigating catastrophic forgetting in continual learning [30] by training tasks sequentially rather than simultaneously.
Pilot study. The aforementioned training details are applied almost identically to the experiments in the pilot study of our main paper. As stated in the main paper, only the classifier was chosen (as a straightforward and practical option) to train on the given Task Arithmetic (or AdaMerging [74]) merged weights for the 8 vision tasks using ViT-B/32. A difference is that, instead of using the test dataset, the training dataset is used in a supervised manner to produce the results closer to the upper bound, in which the difference stands out more. We train the models for only one epoch.
B.2 Baseline Details
We compare our method against a diverse set of baselines, ranging from simple merging strategies to advanced methods leveraging task-specific or shared representations.
Pretrained indicates a model that predicts multiple tasks without additional fine-tuning for task-specific requirements. However, the absence of task-specific information for downstream tasks generally leads to poor performance.
Individual refers to the fine-tuning of individual pre-trained models for each task. Since there is no interference between tasks, it has been regarded as the upper bound of task-specific performance.
Traditional MTL trains a multi-task model by combining the original training data for all tasks. While effective, this approach depends on access to raw data and labels for each task, which is not always practical.
Weight Averaging merges multiple individual models by directly averaging their parameters to create a single model for multi-task learning. Although simple, this method lacks task-specific adjustments.
Fisher Merging [39] uses the Fisher information matrix to estimate the importance of each parameter. Model parameters are merged based on their relative contributions.
RegMean [27] merges models by aligning the inputs of linear layers to minimize the L2 distance between the individual models and the merged model.
Task Arithmetic [23] defines the difference between the fine-tuned and pre-trained model parameters as a task vector. By combining multiple task vectors and adding them to the pre-trained model, it enables multi-task learning.
MagMax [38] merges task vectors [23] by selecting the parameter with the largest magnitude for each position, consolidating knowledge into a single model without retaining task-specific data.
Ties Merging [72] highlight the importance of addressing interference in task arithmetic-based merging. It involves removing redundant parameters from the task vector and resolving parameter sign conflicts.
LiNeS [64] observes that reducing the influence of shallow layers helps prevent distortion of general representations, thereby simplifying merging coefficient selection by allowing them to increase linearly with layer depth.
Consensus TA [63] enhances Task Arithmetic [23] by using Consensus Merging, which retains only weights beneficial to multiple tasks while eliminating irrelevant or task-specific weights. This process uses task-specific binary masks to identify relevant weights and forms a consensus mask to minimize task interference.
Concrete TA [56] utilizes a discrete binary mask to identify a low-dimensional subspace shared across tasks, where Task Arithmetic [23] is applied for model merging.
Concrete AM [56] extends Concrete TA by incorporating the training of the merging coefficients [74] during the merging process.
AdaMerging [74] adaptively learns merging coefficients at the task or layer level by minimizing the entropy of predictions on unlabeled test data.
Representation Surgery [73] reduces representation bias by training a task-specific module that aligns the merged model’s features with those of the individual models.
EMR-Merging [21] involves three steps: Elect, Mask, and Rescale-Merging. These steps select key parameters to form a unified model, apply task-specific masks for each task, and adjust scales to achieve better performance.
WEMoE [58] utilizes a Mixture of Experts (MoE) module to dynamically separate and integrate shared and task-specific knowledge based on input samples. By training the router on unlabeled test data, it optimizes routing weights and improves task-specific flexibility and performance.
Twin-Merging [36] employs the MoE module but introduces a preprocessing step where task-specific knowledge is compressed into sparse vectors via Singular Value Decomposition. During inference, the router integrates shared and task-specific components dynamically based on input.
(a) Fixed Classifier
(b) Classifier Training with Task Arithmetic
(c) Classifier Training with AdaMerging
Figure B: Cross-task evaluations across diverse encoders and heads from different tasks. Each cell $(i,j)$ displays the accuracy when features from task $i$’s visual encoder are classified by task $j$’s classifier. The upper value in a cell shows the absolute task accuracy, and the number below (in parentheses) denotes the relative performance compared to the diagonal. (a) Base (untrained) classifiers generally deteriorate when evaluated on other tasks or using different encoders. (b) Classifiers trained with Task Arithmetic features show mostly reduced performance loss (i.e., higher relative numbers). This suggests that task-specific adjustments effectively refine the task-agnostic representations task-wise. (c) Classifiers trained with AdaMerging features also improve over (a) but perform worse than (b), highlighting the importance of a shared encoder when leveraging task-specific knowledge.
Appendix C More Experimental Results
C.1 Pilot Study
Figure 2(a) shows cross-task evaluations via individual visual encoders with their respective base classifiers; Figure 2(b) shows results from combining individual visual encoders with trained classifiers from Task Arithmetic, while Figure 2(c) presents results using the AdaMerging merged encoder. We observe consistent improvements in cross-task performance when using the classifiers trained on other tasks. For instance, integrating the GTSRB [54] classifier trained on Task Arithmetic features with the SVHN [41] encoder achieves 56.3% accuracy (Figure 2(b)), surpassing the base classifier’s 31.4% (Figure 2(a)). These results suggest that training the classifier on merged features for the GTSRB task also facilitates the alignment of individual SVHN features with the GTSRB task. This demonstrates the trained layer’s enhanced ability to refine task-specific outputs from task-agnostic representations. On the other hand, classifiers trained on AdaMerging features achieve 49.0% accuracy (Figure 2(c)), surpassing the base classifier but performing worse than the classifier trained on Task Arithmetic features. This suggests that even when task-specific layers are trained, the choice of a shared encoder can lead to suboptimal performance. Therefore, it highlights the importance of the shared encoder in effectively leveraging task-specific knowledge.
Method SUN Cars RES. Euro SVH. GTS. MNI. DTD C100 FER Flower Pet PCAM STL Avg.
Pretrained 62.3 59.7 60.7 45.5 31.4 32.6 48.5 43.8 64.2 39.0 66.3 87.4 60.6 97.1 57.1
[…]
Model Tinting
73.3 77.1 94.4 98.8 94.2 98.0 98.8 79.0 84.9 70.6 90.9 91.7 85.4 97.9 88.2
Model Tinting
73.7 78.5 94.2 99.3 94.6 97.8 98.9 80.2 84.4 70.5 92.1 92.2 84.8 98.0 88.5
Table A: Multi-task performance when merging ViT-B/32 models on 14 tasks. $\dagger$ denotes our reproduced results.
Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Avg.
Pretrained 63.8 64.6 65.7 54.5 52.0 43.3 51.7 45.1 55.0
Individual 81.8 86.8 96.9 99.7 97.8 99.1 99.7 82.0 92.9
Weight Averaging 67.7 70.0 75.3 79.5 74.9 60.1 94.4 43.8 70.7
Fisher Merging [39] 68.5 69.9 75.2 80.4 73.2 61.2 94.5 50.7 71.7
RegMean [27] 69.1 71.6 77.6 88.8 83.7 70.2 96.9 54.6 76.6
Task Arithmetic [23] 61.1 65.9 74.0 76.2 88.0 73.9 98.4 53.0 73.8
[…]
Ties Merging [72] 69.1 72.5 80.5 84.0 85.0 71.5 98.1 54.9 77.0
[…]
AdaMerging++ [74] 71.8 80.8 84.1 94.3 91.9 94.5 98.7 69.8 85.7
Ada. w/ Surgery [73] 73.6 81.5 90.4 98.5 93.2 97.4 98.9 77.0 88.8
EMR-Merging [21] 78.6 82.6 95.5 99.2 97.6 98.8 99.6 78.3 91.3
Model Tinting
76.7 87.5 95.4 98.6 96.2 98.7 99.1 82.4 91.8
Model Tinting
77.4 87.0 95.6 98.6 97.0 98.9 99.2 82.7 92.0
Table B: Multi-task performance when merging ViT-B/16 models on 8 tasks.
Figure C: Multi-task performance when merging RoBERTa models on 8 tasks. Yellow indicates training only coefficients, green indicates training only the classifier, and blue indicates training both in our method.
C.2 NLP Tasks
We conduct experiments to investigate how the training of the classifier and merging coefficients, respectively, affects the performance of NLP tasks in Model Tinting , as presented in Table 3 of the main paper. Figure C shows the results of merging RoBERTa [35] models for 8 NLP tasks using the predictions of individual models as guidance, optimized through the cross-entropy loss. As mentioned in the main paper, evaluation metrics differ across tasks: the Matthews correlation coefficient is used for CoLA [65], the average of Pearson and Spearman correlations is applied to STS-B [5], and the accuracy is used for all other tasks [53, 12, 24, 66, 49, 14]. The approach where both the classifier and coefficients are trained simultaneously corresponds to Model Tinting . To examine the role of classifier training in NLP tasks, ablation experiments are conducted by removing specific components. The results show that training only the coefficients leads to the lowest performance while training only the classifier achieves a relatively high performance. Furthermore, training both the classifier and coefficients together demonstrated a complementary effect, achieving the highest performance.
C.3 Vision Tasks
Table A presents the results of merging ViT-B/32 across 14 tasks. Our method outperforms the current state-of-the-art, EMR-Merging [21], by an average of 2.4%p. Notably, Model Tinting outperforms individual models on datasets such as Cars [31], DTD [7], Flowers [42], Pets [45], and STL10 [9], demonstrating its robustness and effectiveness across a wide range of tasks and datasets, regardless of the number of tasks. Table B shows the results of merging ViT-B/16 across 8 vision tasks. While other methods exhibit significant performance gaps compared to individual models, particularly on datasets like Cars and DTD, Model Tinting consistently achieves performance close to that of individual models.
Layer SUN Cars RES. Euro SVH. GTS. MNI. DTD Avg.
1 73.5 76.8 95.2 99.7 95.9 98.4 99.4 79.4 89.8
2 73.5 76.8 95.6 99.9 96.2 98.6 99.4 79.4 89.9
3 73.9 76.8 95.4 99.8 96.2 98.6 99.3 79.5 89.9
4 73.8 76.7 95.4 99.9 96.5 99.0 99.4 79.4 90.0
5 74.1 77.0 95.5 99.8 96.5 98.6 99.4 79.4 90.0
6 73.9 76.7 95.6 99.9 96.3 98.3 99.4 79.4 89.9
7 74.1 77.1 95.5 99.7 96.4 98.8 99.5 79.4 90.1
8 74.1 75.6 95.3 99.8 95.9 98.6 99.6 79.4 89.8
9 74.1 77.3 95.6 99.9 96.1 98.3 99.2 79.4 90.0
10 74.5 77.3 95.7 99.7 96.0 98.4 99.4 79.4 90.0
11 74.5 77.2 95.5 99.8 96.3 98.5 99.5 79.4 90.1
12 74.3 76.7 94.7 99.8 95.8 98.2 99.3 79.4 89.8
Classifier 73.7 77.6 94.7 99.0 95.4 98.2 99.1 79.1 89.6
Late 69.6 73.1 94.2 99.7 96.0 98.6 98.4 79.1 88.6
Early 70.0 75.2 93.4 99.3 93.9 98.1 99.2 79.3 88.5
All 67.1 69.5 94.2 98.4 95.1 97.7 98.8 75.4 87.0
Table C: Performance details of merged ViT-B/32 models when training different layers with merging coefficients for task-specific adjustments. We observe that consistently high performance can be achieved regardless of which task-specific single layer is trained.
Appendix D More Empirical Analyses
D.1 Detailed Results on Task-specific Layers
Our proposed method, Model Tinting , introduces the concept of incorporating a task-specific layer to balance shared and task-specific representations during model merging. This approach allows effective adaptation while maintaining efficiency to address task conflicts in multi-task learning. ViT-B/32 is employed and initialized using Task Arithmetic-merged weights across all layers. When a particular layer is marked as task-specific, it is split into unique versions for each task, resulting in eight task-specific layers. While these layers are trained exclusively on corresponding task data, the remaining layers update only the merging coefficients using data from all tasks. Table C presents detailed task-level results. ‘Early’ refer to layers 1 to 6-th, while ‘Late’ refer to layers 7 to 12-th. Training both the merging coefficients and task-specific layers achieves performance comparable to or slightly better than training merging coefficients alongside the classifier. These findings highlight the minimal impact of task-specific layer choice on overall performance, reaffirming the robustness of our approach. The results also show consistent average performance across layers, demonstrating that merging coefficients effectively balance representations across tasks, mitigating biases even when other layers remain frozen.
D.2 On Loss Functions
Figure D: Comparison of training with various loss functions. The yellow bars represent training only the coefficients, while the blue bars indicate the joint training of the classifier and coefficients. The dashed line corresponds to training with ground truth labels using cross-entropy loss, which is the upper bound.
(a) Before training (left: EM Merging   right: GT)
(b) After training (left: EM Merging   right: GT)
Figure E: t-SNE visualization results on the DTD dataset before (a) and after (b) entropy minimization training. The plots on the left use merged model predictions, while those on the right use ground truth labels. Training improves clustering, but prediction-ground truth mismatch remains.
In Figure D, we analyze the impact of various loss functions on aligning predictions between individual models and the merged model. When training only the coefficients, distance-based losses (e.g., L1, Smooth-L1, L2, and Cosine) demonstrate relatively lower performance compared to distribution-based losses (e.g., KL-divergence, JS-divergence, cross-entropy, and entropy). In contrast, when both the classifier and coefficients are trained jointly, most loss functions achieve strong performance under our method. However, entropy minimization, as explored in previous works [58, 74], leads to a significant decline in performance during joint training.
Notably, in Table 2 of the main paper, we observe that the entropy loss often fails to provide gradients in the correct direction due to its divergence from the ground truth (GT)-based loss. The table highlights that the cross-entropy loss, in contrast, aligns more closely with the GT, resulting in more accurate gradient directions. This highlights the importance of incorporating label information for guiding classifier training in the merged model. To further support this point, Figure E visualizes the t-SNE visualization results for the DTD dataset before and after entropy minimization training. While feature clustering improves after training, as shown in (b) compared to (a), a noticeable discrepancy remains between the predictions and the ground truth labels. These observations emphasize the importance of label guidance, even if incomplete, for achieving better alignment with the GT and improving the performance of the merged model.
SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD
Entropy Before 0.73 0.82 0.84 0.86 0.96 0.91 1.00 0.61
After 0.38 0.31 0.94 0.92 0.88 0.96 0.98 0.38
$\Delta$ (-0.35) (-0.51) (+0.1) (+0.06) (-0.08) (+0.05) (-0.02) (-0.23)
Ours Before 0.82 0.92 0.98 1.00 0.99 1.00 1.00 0.85
After 0.86 0.95 1.00 1.00 0.99 1.00 1.00 0.90
$\Delta$ (+0.04) (+0.03) (+0.02) (+0.00) (+0.00) (+0.00) (+0.00) (+0.05)
(a) ViT-B/16
SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD
Entropy Before 0.85 0.93 0.95 0.98 0.96 0.96 1.00 0.79
After 0.63 0.76 0.94 0.81 0.79 0.95 0.82 0.83
$\Delta$ (-0.22) (-0.17) (-0.01) (-0.17) (-0.17) (-0.01) (-0.18) (+0.04)
Ours Before 0.87 0.96 0.98 1.00 0.98 1.00 1.00 0.89
After 0.89 0.97 1.00 1.00 0.99 1.00 1.00 0.91
$\Delta$ (+0.02) (+0.01) (+0.02) (+0.00) (+0.01) (+0.00) (+0.00) (+0.02)
(b) ViT-L/14
Table D: Spearman correlation of losses with ground truth cross-entropy loss for (a) ViT-B/16 and (b) ViT-L/14. Values closer to 1 indicate that the corresponding loss function exhibits better alignment with the loss computed using the ground truth.
Loss correlation. We conduct further experiments to validate whether the loss correlation shown in Table 2 of the main paper remains consistent across different backbones (ViT-{B/16, L/14}). Table D shows the loss correlation results for 8 vision tasks using both the merged ViT-B/16 and ViT-L/14 models. When both the merging coefficients and the classifier are trained simultaneously, we observe a significant drop in correlation after training with entropy minimization. In contrast, the cross-entropy loss employed in our self-labeling approach maintains consistently high correlation across tasks. These findings are consistent with the results reported for ViT-B/32 in the main paper, suggesting that our self-labeling approach with the cross-entropy loss could be more effective than the entropy loss when training the classifier alongside the coefficients.
(a) Comparison of average accuracy for ViT-B/32, ViT-B/16, and ViT-L/14 across training iterations.
(b) Average accuracy of ViT-B/32 across training iterations under different learning rates.
Figure F: Learning curves of our method on the 8 vision tasks across several models and learning rates. Our method demonstrates robust and fast convergence, regardless of model size and learning rate.
D.3 Convergence Analysis
Figure F illustrates the learning curve of Model Tinting applied to the 8 vision tasks over training iterations. In Figure F, we compare the performance of ViT-B/32, ViT-B/16, and ViT-L/14 models. The results demonstrate that our method achieves near-optimal performance within the first 100 iterations and converges quickly, regardless of model size. Notably, larger models like ViT-L/14 consistently achieve higher final accuracies, followed by ViT-B/16 and ViT-B/32, which highlights the advantages of model capacity in capturing task-specific knowledge more effectively. Figure F explores the impact of learning rates on the convergence of ViT-B/32. Except for the case with the lowest learning rate, the figure shows that our method converges quickly to a similarly high level of performance across most learning rates. These observations demonstrate the robustness and efficiency of Model Tinting across different model scales and hyperparameter settings.
D.4 Sparsity Visualization
The main paper highlights that sparsity in merging coefficients plays a critical role in reducing task conflicts during multi-task model merging. To confirm this trend across various backbones, Figure GH and I provide the layer-wise merging coefficients for ViT-B/32, ViT-B/16 and ViT-L/14. All models are optimized using the cross-entropy loss with the predictions of individual models as guidance. In these figures, (a) represents training only the coefficients, while (b) includes joint training of the classifier and coefficients. Joint training consistently increases the proportion of near-zero coefficients (from 35.3% to 56.6% for ViT-B/16 and from 23.3% to 55.0% for ViT-L/14). Correspondingly, the average accuracy also improves (from 86.8% to 91.8% for ViT-B/16 and from 91.5% to 93.8% for ViT-L/14). These results indicate that training the classifier complements sparsity in merging coefficients, effectively reducing task conflicts and enhancing performance.
(a) Training coefficients only
(b) Joint training
Figure G: Impact of joint training on merging coefficient sparsity on ViT-B/32.
(a) Training coefficients only
(b) Joint training
Figure H: Impact of joint training on merging coefficient sparsity on ViT-B/16.
(a) Training coefficients only
(b) Joint training
Figure I: Impact of joint training on merging coefficient sparsity on ViT-L/14.